P=C/(1+ae^kt) can you solve this for t?
Answer by stanbon(75887)

P = C/(1+ae^kt) can you solve this for t?
------------------
(1+ae^(kt)) = P/C
ae^(kt) = (P/C)-1
---
e^(kt) = [(p-C)/aC]
---
kt = ln[(p-C)/aC]
t = [ln[(p-C)/aC]]/k
=========================
Cheers,
Stan H.
=========================